# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ Appendix A Appendix: PXP Model
    where $\Delta_{l,\pm}=\left(\Sigma_{(l,1)}\pm\Sigma_{(l,0)}\right)/2$ and we used $\gamma_{\mathrm{eff}}=2\gamma$. Besides, we set the coefficients $\Theta_{2}=\sum_{s}\Sigma_{(1,s)}$, $\Theta_{1}=\sum_{s>s^{\prime}}\left(\Sigma_{(1,s)}\Sigma_{(1,s^{\prime})}\right)+736\gamma^{2}/25$, and $\Theta_{0}=\Sigma_{(1,0)}\Sigma_{(1,1)}\Sigma_{(1,2)}+\left(576\Sigma_{(1,0)}+32\Sigma_{(1,2)}\right)\gamma^{2}/5$.
In $\gamma=0$ case, all of the first-order eigenvalues $\lambda_{l}^{(1)}$ are purely imaginary. As the dephasing rate $\gamma$ increases and exceeds the $\mathbb{PT}$-symmetry breaking point, only one of $\lambda_{l=l^{\prime\prime}}^{(1)}$ in each $l^{\prime\prime}$ layer always remain imaginary, while other $\lambda_{l=l^{\prime\prime}}^{(1)}$ will become complex. On the other hand, the real part of $\lambda_{l=l^{\prime}}^{(1)}$ becomes nonzero once $\gamma>0$, which explains why $\mathbb{PT}$-symmetry breaking does not exist in the $l^{\prime}$ layer.
Furthermore, for $J_{\mathrm{x},j}=0.1$, the parameters $\langle E_{0}|\hat{\Sigma}|E_{0}\rangle\approx 0.064$, $\langle E_{1}|\hat{\Sigma}|E_{1}\rangle\approx 0.028$, $\langle E_{2}|\hat{\Sigma}|E_{2}\rangle\approx 0.0061$ can be calculated by using the non-equilibrium Green’s function method in Eq. (S8). The first-order eigenvalue as the dephasing rate increases under these parameters were shown in the main text, illustrating the LPTS breaking point approximately aligns with the numerical results.
Although the $|l|=N$ momentum layer is unrelated to LPTS breaking, for completeness we also derive the perturbative correction for this layer. In this case, we simply use non-degenerate perturbation theory and obtain $\lambda^{\prime}_{N}\approx\lambda^{(0)}_{N}+\lambda^{(1)}_{N}$, with the first-order eigenvalue given by
$\lambda^{(1)}_{|l|=N}=\pm\Big{\langle}\!\!\Big{\langle}\lambda_{(N,0)}^{(0)}\Big{|}\left(\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}+\mathcal{E}\right)\Big{|}\lambda_{(N,0)}^{(0)}\Big{\rangle}\!\!\Big{\rangle}$ =±$=\pm\gamma_{\mathrm{eff}}\sum_{j}^{N}\big{(}\langle E_{N}|\hat{\sigma}_{j}^{z}|E_{N}\rangle\times\langle E_{0}|\hat{\sigma}_{j}^{z}|E_{0}\rangle\big{)}\pm i\left(\langle E_{N}|\hat{\Sigma}|E_{N}\rangle-\langle E_{0}|\hat{\Sigma}|E_{0}\rangle\right)$ (S20)
=±$=\pm\gamma_{\mathrm{eff}}\sum_{j}^{N}\langle+|^{\otimes N}\hat{\sigma}_{j}^{z}|+\rangle^{\otimes N}\langle-|^{\otimes N}\hat{\sigma}_{j}^{z}|-\rangle^{\otimes N}\pm i\left(\langle E_{N}|\hat{\Sigma}|E_{N}\rangle-\langle E_{0}|\hat{\Sigma}|E_{0}\rangle\right)$
=±$=\pm i\left(\langle E_{N}|\hat{\Sigma}|E_{N}\rangle-\langle E_{0}|\hat{\Sigma}|E_{0}\rangle\right).$
@@
Figure S1: (a) Schematic of a general open quantum many-body system with approximate scars. The scar and thermal subsystems lose their quantum coherence to the bath with the dephasing rate $\gamma$, while interacting via the coupling strength $V$. (b) The Liouvillean spectral transition originates from the competition between self-energy, $\mathcal{E}$, and dissipation $\mathcal{D}$. A set of eigenmodes in Liouville space, denoted by blue concentric circles, are surrounded by the self-energy shell $\mathcal{E}$, due to the thermal subspace. The latter suppresses those dephasing processes (wavy arrows) that do not exceed the boundary of the self-energy shell (yellow circle).
Appendix S2 A3. Scaling behavior
@@
From the perspective of perturbation theory, the LPTS breaking phenomenon originates from the competition between dissipation $\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}$ and the self-energy $\mathcal{E}$, as summarized in Fig. S1. We can estimate the typical LPTS breaking point:
@@
where $\|\hat{A}\|=\sqrt{\lambda_{\mathrm{max}}(\hat{A}^{\dagger}\hat{A})}$ refers to the matrix norm of $\hat{A}$. The dissipation superoperator scales with $N$, $\|\mathcal{D}^{\prime}\|=\sum_{j=1}^{N}\|\hat{\sigma}_{j}^{z}\otimes\hat{\sigma}_{j}^{z}\|\sim N$. To estimate the scaling of self-energy for large $N$, recall the self-consistency condition, Eq. (S8), and focus on the SCL model, with $\hat{S}_{0}$ in Eq. (S4). The maximal energy is $NJ$, so we have $\|E\hat{I}-\hat{S}_{0}\|$ scaling as $N$. According to Eq. (S9),
$\big{\|}\hat{\Gamma}\big{\|}=\sqrt{\max{\left(\sum_{j=1}^{N-1}J_{\mathrm{x},j}\hat{P}^{j,j+1}_{p_{\mathrm{S}}}\right)}}.$ (S22)
The possible value of $\hat{P}^{j,j+1}_{p_{\mathrm{S}}}$ is 00 and $1$, so $\|\hat{\Gamma}\|=\sqrt{\sum_{j=1}^{N-1}J_{\mathrm{x},j}}\sim\sqrt{N}$, where we assume that all the couplings $J_{\mathrm{x},j}$ are at the same order of magnitude. Approximately, from the self-consistence equation, we assume that $\|\hat{\Sigma}\|$ increase slower than $\|E\hat{I}-\hat{S}_{0}\|$, and we have

### v3

@@ Abstract
    where $\Delta_{l,\pm}=\left(\Sigma_{(l,1)}\pm\Sigma_{(l,0)}\right)/2$ and we used $\gamma_{\mathrm{eff}}=2\gamma$ due to two neighboring sites (dimer) are used to define a ‘spin-$1/2$’ in the scar subspace (i.e., $N=L/2$). Besides, we set the coefficients $\Theta_{2}=\sum_{s}\Sigma_{(1,s)}$, $\Theta_{1}=\sum_{s>s^{\prime}}\left(\Sigma_{(1,s)}\Sigma_{(1,s^{\prime})}\right)+736\gamma^{2}/25$, and $\Theta_{0}=\Sigma_{(1,0)}\Sigma_{(1,1)}\Sigma_{(1,2)}+\left(576\Sigma_{(1,0)}+32\Sigma_{(1,2)}\right)\gamma^{2}/5$.
Furthermore, for $J_{\mathrm{x},j}=0.1$, the parameters $\langle E_{0}^{\mathrm{S}}|\hat{\Sigma}|E_{0}^{\mathrm{S}}\rangle\approx 0.064$, $\langle E_{1}^{\mathrm{S}}|\hat{\Sigma}|E_{1}^{\mathrm{S}}\rangle\approx 0.028$, $\langle E_{2}^{\mathrm{S}}|\hat{\Sigma}|E_{2}^{\mathrm{S}}\rangle\approx 0.0061$ can be calculated by using the non-equilibrium Green’s function method in Eq. (S13). The first-order eigenvalue as the dephasing rate increases under these parameters were shown in the main text, illustrating the spectral transition point approximately aligns with the numerical results.
For the case $|l|=N$, we simply use non-degenerate perturbation theory and obtain $\lambda^{\prime}_{N}\approx\lambda^{(0)}_{N}+\lambda^{(1)}_{N}$, with the first-order eigenvalue given by
$\lambda^{(1)}_{|l|=N}=\pm\Big\langle\!\!\Big\langle\lambda_{(N,0)}^{(0)}\Big|\left(\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}+\mathcal{E}\right)\Big|\lambda_{(N,0)}^{(0)}\Big\rangle\!\!\Big\rangle$ $=\pm\gamma_{\mathrm{eff}}\sum_{j}^{N}\big(\langle E_{N}^{\mathrm{S}}|\hat{\mathbb{S}}_{j}^{z}|E_{N}^{\mathrm{S}}\rangle\times\langle E_{0}^{\mathrm{S}}|\hat{\mathbb{S}}_{j}^{z}|E_{0}^{\mathrm{S}}\rangle\big)\pm i\left(\langle E_{N}^{\mathrm{S}}|\hat{\Sigma}|E_{N}^{\mathrm{S}}\rangle-\langle E_{0}^{\mathrm{S}}|\hat{\Sigma}|E_{0}^{\mathrm{S}}\rangle\right)$ (S20)
$=\pm\gamma_{\mathrm{eff}}\sum_{j}^{N}\langle+|^{\otimes N}\hat{\mathbb{S}}_{j}^{z}|+\rangle^{\otimes N}\langle-|^{\otimes N}\hat{\mathbb{S}}_{j}^{z}|-\rangle^{\otimes N}\pm i\left(\langle E_{N}^{\mathrm{S}}|\hat{\Sigma}|E_{N}^{\mathrm{S}}\rangle-\langle E_{0}^{\mathrm{S}}|\hat{\Sigma}|E_{0}^{\mathrm{S}}\rangle\right)$
$=\pm i\left(\langle E_{N}^{\mathrm{S}}|\hat{\Sigma}|E_{N}^{\mathrm{S}}\rangle-\langle E_{0}^{\mathrm{S}}|\hat{\Sigma}|E_{0}^{\mathrm{S}}\rangle\right).$
@@
A few comments are in order. The layer indices can be split into two groups, where QMBS layers $l=-N,-N+2,\cdots,N$ represent layers undergoing spectral transition in Fig. 1 of the main text, and $l=-N+1,-N+3,\cdots,N-1$ are the remaining layers without spectral transition. The layers $l=\pm N$ are special in that they contain a single level $s=0$. Consequently, as we show in above, these QMBS Liouvillian eigenvalues remain on the imaginary axis for arbitrary values of the dephasing, while the other $l$ sectors generally distribute across the complex plane. Second, we note that our labeling of QMBS layers is unrelated with inversion symmetry quantum number (see Sec. B for a discussion of symmetries of the Liouvillian). The latter is given by $p=(-1)^{l}$, therefore the spectral transition is found in different inversion-symmetric subspaces depending on the parity of $N$.
S1.2 A2. The cases without $\mathcal{T}_{-}$ symmetry: PXP model
The scar subspace of the PXP model can also be approximately described using the Hamiltonian decomposition in Eq. (S1). However, the QMBS Hamiltonian $\hat{S}_{0}$ and self-energy $\hat{\Sigma}$ should be replaced by an effective spin-1 paramagnet in the dimerized picture [24, 97]. First, we use an artificial $S=1$ “blockspin” by defining $|0\rangle_{b}:=|\downarrow\!\downarrow\rangle_{b_{\alpha},b_{\beta}},|+\rangle_{b}:=|\downarrow\uparrow\rangle_{b_{\alpha},b_{\beta}},|-\rangle_{b}:=|\uparrow\downarrow\rangle_{b_{\alpha},b_{\beta}}$, where $b_{\alpha}(b_{\beta})$ represents the left(right) site in the $b$-th block. Letting $N=L/2$ be the number of the blocks, the QMBS Hamiltonian $\hat{S}_{0}$ and self-energy $\hat{\Sigma}$ can be written as
$\hat{S}_{0}=\sqrt{2}\sum_{b=1}^{N}\hat{\tau}_{b}^{x},\quad\quad\hat{\Sigma}=-\sum_{b=1}^{N}(|+,0\rangle+|0,-\rangle)\left\langle+,-\right|_{b,b+1}+\mathrm{h.c.},$ (S21)
where $\hat{\tau}_{b}$ is the standard spin-$1$ Pauli matrices for $b$-th block in the PXP model. In this picture, the initial state and the eigenstates can be expressed as
$\left|\mathbb{Z}_{2}\right\rangle=\bigotimes^{N}_{b=1}|\downarrow\uparrow\rangle_{b_{\alpha},b_{\beta}}=\bigotimes^{N}_{b=1}|+\rangle_{b},\quad|E_{s}^{\mathrm{S}}\rangle=\sqrt{\frac{1}{\mathcal{N}_{s}}}(\hat{J}^{-})^{N-s}\bigotimes^{N}_{b=1}|\widetilde{+}\rangle_{b},$ (S22)
where $\hat{J}^{\pm}=\mp i\sum_{b}\hat{\tau}^{\pm}_{b}$ is the raising (lowering) operator, with $\hat{\tau}^{\pm}_{b}=\hat{\tau}^{y}_{b}\pm i\hat{\tau}^{z}_{b}$. The normalization factor $\mathcal{N}_{s}$ reads
$\mathcal{N}_{s}=\prod_{M=s-N+1}^{N}\left[(N\left(N+1\right)-M(M-1)\right],$ (S23)
where the index is chosen from $s=0,1,2,\cdots,L$. In addition, $|\widetilde{\pm}\rangle_{b}$ and $|\widetilde{0}\rangle_{b}$ are the eigenvectors of the local operator $\hat{\tau}_{b}^{x}$ with positive (negative) and zero eigenvalue.
Considering the Rydberg projector $P_{\mathrm{Ryd}}=\prod_{b}(\hat{I}-|+,-\rangle\langle+,-\mid)_{b,b+1}$, the eigenstates of PXP model can be approximated by $|E^{\mathrm{S}\ast}_{s}\rangle=P_{\mathrm{Ryd}}|E_{s}^{\mathrm{S}}\rangle$. Based on first-order perturbation theory, the energy shift $\Delta E_{s}=E_{s}^{\mathrm{S}\ast}-E_{s}^{\mathrm{S}}$ is given as
$\Delta E_{s}={\left\langle E_{s}^{\mathrm{S}}\right|\hat{\Sigma}\left|E_{s}^{\mathrm{S}}\right\rangle}=-\frac{1}{\mathcal{N}_{s}}\sum_{b=1}^{N}\left\langle E_{s}^{\mathrm{S}}\right|(|+,0\rangle+|0,-\rangle)\left\langle+,-\right|_{b,b+1}+\mathrm{h.c.}\mid E_{s}^{\mathrm{S}}\rangle.$ (S24)
After some algebraic manipulations, the energy correction from all the sites is thus estimated as [24]
${\left\langle E_{s}^{\mathrm{S}}\right|\hat{\Sigma}\left|E_{s}^{\mathrm{S}}\right\rangle}=-\frac{\sqrt{2}}{8}N\frac{\left|c_{+}\right|^{2}-\left|c_{-}\right|^{2}}{\mathcal{N}_{s}},$ (S25)
where the coefficient $c_{\pm}$ reads
$c_{+}$ $=\prod_{M=s-N-1}^{N-2}\sqrt{\left(N-2\right)\left(N-1\right)-M(M-1)},$ (S26)
$c_{-}$ $=s(s-1)(s-2)(s-3)\prod_{M=s-N+3}^{N-2}\sqrt{\left(N-2\right)\left(N-1\right)-M(M-1)}.$
Then we put the results of Eq. (S25) into Eq. (S15) and obtain the diagonal matrix $\mathcal{E}_{l}$.
The next step is to calculate the projection of $\mathcal{D}^{\prime}$ to the $l$th layer, $\mathcal{D}_{l}^{\prime}$. In the spin-$1$ picture, the dephasing term is expressed as
$\mathcal{D}^{\prime}=\sum_{b}\hat{D}_{b}^{\alpha}\otimes\hat{D}_{b}^{\alpha}+\sum_{b}\hat{D}_{b}^{\beta}\otimes\hat{D}_{b}^{\beta},$ (S27)
$\hat{D}_{b}^{\alpha}=-|\downarrow\!\uparrow\rangle\langle\downarrow\!\uparrow|_{b_{\alpha},b_{\beta}}-|\downarrow\!\downarrow\rangle\langle\downarrow\!\downarrow|_{b_{\alpha},b_{\beta}}+|\uparrow\!\downarrow\rangle\langle\uparrow\!\downarrow|_{b_{\alpha},b_{\beta}}=-\hat{\tau}_{b}^{z}-|0\rangle\langle 0|_{b},$
$\hat{D}_{b}^{\beta}=|\downarrow\!\uparrow\rangle\langle\downarrow\!\uparrow|_{b_{\alpha},b_{\beta}}-|\downarrow\!\downarrow\rangle\langle\downarrow\!\downarrow|_{b_{\alpha},b_{\beta}}-|\uparrow\!\downarrow\rangle\langle\uparrow\!\downarrow|_{b_{\alpha},b_{\beta}}=\hat{\tau}_{b}^{z}-|0\rangle\left\langle 0\right|_{b}.$
Because of the Rydberg projector, the dissipator term is not exactly the linear Zeeman term as in the CL model. If we only consider the linear Zeeman term $\hat{\tau}^{z}$, we obtain something similar to the CL model in that the dissipator matrix $\mathcal{D}_{l}^{\prime}$ couples only $|\lambda_{(l,s)}^{(0)}\rangle\!\rangle$ and $|\lambda_{(l,s\pm 1)}^{(0)}\rangle\!\rangle$,
$\mathcal{D}_{l}^{\prime}=\sum_{s=0}^{L-|l|}d_{(l,s)}^{(0)}\ \big|\lambda_{(l,s)}^{(0)}\big\rangle\!\big\rangle\big\langle\!\big\langle\lambda_{(l,s)}^{(0)}\big|+\sum_{k=1}^{L-|l|}\sum_{s=0}^{L-|l|-k}d_{(l,s)}^{(k)}\Big[\ \big|\lambda_{(l,s)}^{(0)}\big\rangle\!\big\rangle\big\langle\!\big\langle\lambda_{(l,s+k)}^{(0)}\big|+\mathrm{h.c.}\Big],$ (S28)
where $d^{(k)}_{(l,s)}$ represents the coupling coefficient between the $s$-th and $(s+k)$-th degenerate eigenstates in the $l$-momentum layer. It follows that $d^{(k)}_{(l,s)}=0$ when $k\neq 1$. For PXP model, the corresponding eigenmodes and eigenvalues are written as
$|\lambda_{(l,s)}^{(0)}\rangle\!\rangle=|E_{l+s}^{\mathrm{S}}\rangle\otimes|E_{s}^{\mathrm{S}}\rangle,\quad\quad\quad\quad\lambda_{l}^{(0)}=i(E_{l+s}^{\mathrm{S}}-E_{s}^{\mathrm{S}})=2\sqrt{2}ilJ.$ (S29)
Using the translation symmetry of $\sum_{i}\hat{\tau}^{z}_{i}$, we have
$d_{(l,s)}^{(1)}$ $=2\sum_{b=1}^{N}\Big[\langle E_{s}^{\mathrm{S}}|\hat{\tau}^{z}_{b}|E_{s+1}^{\mathrm{S}}\rangle\langle E_{s+|l|}^{\mathrm{S}}|\hat{\tau}^{z}_{b}|E_{s+|l|+1}^{\mathrm{S}}\rangle\Big]=\frac{2}{N}\bigg[\Big\langle E_{s}^{\mathrm{S}}\Big|\sum_{b}\hat{\tau}^{z}_{b}\Big|E_{s+1}^{\mathrm{S}}\Big\rangle\Big\langle E_{s+|l|}^{\mathrm{S}}\Big|\sum_{b}\hat{\tau}^{z}_{b}\Big|E_{s+|l|+1}^{\mathrm{S}}\Big\rangle\bigg].$ (S30)
Noting that $\sum_{b}\hat{\tau}^{z}_{b}=(1/2)(\hat{J^{+}}+\hat{J^{-}})$ and using the definition of $|E_{s}^{\mathrm{S}}\rangle$ in the Eq. (S22), we obtain
$\Big\langle E_{s}^{\mathrm{S}}\Big|\sum_{b}\hat{\tau}^{z}_{b}\Big|E_{s+1}^{\mathrm{S}}\Big\rangle=$ $\frac{1}{2}\left\langle E_{s}^{\mathrm{S}}\left|\hat{J}^{+}\right|E_{s+1}^{\mathrm{S}}\right\rangle+\frac{1}{2}\left\langle E_{s}^{\mathrm{S}}\left|\hat{J}^{-}\right|E_{s+1}^{\mathrm{S}}\right\rangle$ (S31)
$=$ $\frac{1}{2}\sqrt{\frac{\mathcal{N}_{s}}{\mathcal{N}_{s+1}}}\left(\bigotimes^{N}_{b=1}\langle\widetilde{+}|_{b}\right)\left(\hat{J}^{-}\right)^{s-(s+1)}\hat{J}^{+}\left(\bigotimes^{N}_{b=1}|\widetilde{+}\rangle_{b}\right)+$
$\frac{1}{2}\sqrt{\frac{\mathcal{N}_{s}}{\mathcal{N}_{s+1}}}\left(\bigotimes^{N}_{b=1}\langle\widetilde{+}|_{b}\right)\left(\hat{J}^{-}\right)^{s-(s+1)}\hat{J}^{-}\left(\bigotimes^{N}_{b=1}|\widetilde{+}\rangle_{b}\right)$
$=$ $0+\frac{1}{2}\sqrt{\frac{\mathcal{N}_{s}}{\mathcal{N}_{s+1}}}.$
Hence, the coefficient $d_{(l,s)}^{(1)}$ can be written in a compact form
$d_{(l,s)}^{(1)}=\frac{1}{2N}\sqrt{\frac{\mathcal{N}_{s}}{\mathcal{N}_{s+1}}}\sqrt{\frac{\mathcal{N}_{s+|l|}}{\mathcal{N}_{s+|l|+1}}}=\frac{1}{2N}$ $\sqrt{N\left(N+1\right)-(N-s)(N-s-1)}$ (S32)
$\times$ $\sqrt{N\left(N+1\right)-(N-s-|l|)(N-s-|l|-1)}.$
[…]
$d_{(l,s)}^{(0)}$ $=\frac{1}{2}\sum_{b=1}^{N}\Big[\langle E_{s}^{\mathrm{S}}|(\hat{\tau}^{x}_{b})^{2}|E_{s}^{\mathrm{S}}\rangle\langle E_{s+|l|}^{\mathrm{S}}|(\hat{\tau}^{x}_{b})^{2}|E_{s+|l|}^{\mathrm{S}}\rangle\Big]=\frac{N}{2}\langle E_{s}^{\mathrm{S}}|(\hat{\tau}^{x}_{1})^{2}|E_{s}^{\mathrm{S}}\rangle\langle E_{s+|l|}^{\mathrm{S}}|(\hat{\tau}^{x}_{1})^{2}|E_{s+|l|}^{\mathrm{S}}\rangle.$ (S33)
$d_{(l,s)}^{(2)}$ $=\frac{1}{8}\sum_{b=1}^{N}\Big[\langle E_{s}^{\mathrm{S}}|(\hat{\tau}^{-}_{b})^{2}|E_{s+2}^{\mathrm{S}}\rangle\langle E_{s+|l|}^{\mathrm{S}}|(\hat{\tau}^{-}_{b})^{2}|E_{s+|l|+2}^{\mathrm{S}}\rangle\Big]=\frac{N}{8}\langle E_{s}^{\mathrm{S}}|(\hat{\tau}^{-}_{1})^{2}|E_{s+2}^{\mathrm{S}}\rangle\langle E_{s+|l|}^{\mathrm{S}}|(\hat{\tau}^{-}_{1})^{2}|E_{s+|l|+2}^{\mathrm{S}}\rangle.$
We first define the state $|T_{N,M}\rangle\equiv|E_{s}^{\mathrm{S}}\rangle$ with total spin $N$ and total magnetic quantum number M=sNM=s-N, then decompose it via Clebsch-Gordan coefficients
$|T_{N,M}\rangle=\sum_{m_{1}}C_{1,m_{1};N-1,M-m_{1}}^{N,M}|T_{1,m_{1}}\rangle\otimes|T_{N-1,M-m_{1}}\rangle,$ (S34)
where $m_{1}$ is magnetic quantum number of the first block. The probability for the first block to have $m_{1}=\pm 1$ is:
$P_{\pm}=\left|C_{1,\pm 1;N-1,M\mp 1}^{N,M}\right|^{2}=\frac{(N\pm M)(N\pm M-1)}{2N(2N-1)}.$ (S35)
The fully symmetric state $|T_{N,M}\rangle$ ensures identical expectation values for any single-block operator. Thus, $\langle(\hat{\tau}^{x}_{b})^{2}\rangle$ is block-independent. Since $(\hat{\tau}^{x}_{1})^{2}$ has eigenvalue 1 for $m_{1}=\pm 1$ and 0 for $m_{1}=0$, the total probability gives:
$\langle T_{N,M}|(\hat{\tau}^{x}_{1})^{2}|T_{N,M}\rangle=P_{+}+P_{-}=\frac{(N+M)(N+M-1)+(N-M)(N-M-1)}{2N(2N-1)}=\frac{N^{2}+M^{2}-N}{N(2N-1)}.$ (S36)
In addition, $(\hat{\tau}^{-}_{1})^{2}$ reduces $m_{1}$ by 2 units, mapping $|T_{N,M+2}\rangle$ to a state with total $\tau^{x}=M$. Only the $m_{1}=+1$ component contributes:
$(\hat{\tau}^{-}_{1})^{2}|T_{N,M+2}\rangle=2C_{1,+1;N-1,M+1}^{N,M+2}|T_{1,-1}\rangle\otimes|T_{N-1,M+1}\rangle.$ (S37)
where the coefficients for maximal spin coupling are:
$C_{1,+1;N-1,M+1}^{N,M+2}=\sqrt{\frac{(N+M+2)(N+M+1)}{2N(2N-1)}},\quad\quad C_{1,-1;N-1,M+1}^{N,M}=\sqrt{\frac{(N-M)(N-M-1)}{2N(2N-1)}}.$ (S38)
Then the inner product gives:
$\langle T_{N,M}|(\hat{\tau}^{-}_{1})^{2}|T_{N,M+2}\rangle=2C_{1,+1;N-1,M+1}^{N,M+2}C_{1,-1;N-1,M+1}^{N,M}=\frac{\sqrt{(N+M+2)(N+M+1)(N-M)(N-M-1)}}{N(2N-1)}.$ (S39)
Using the results in Eq. (S33), Eq. (S36) and Eq. (S39) and substituting $M$ with sNs-N, the coefficient $d^{(k)}_{(l,s)}$ of PXP model can be concluded as
$d^{(0)}_{(l,s)}$ $=\frac{1}{2N(2N-1)^{2}}\Big[N(N-1)+(N-s)^{2}\Big]\Big[N(N-1)+(N-s-|l|)^{2}\Big],$ (S40)
$d^{(1)}_{(l,s)}$ $=\frac{1}{2N}\sqrt{N\left(N+1\right)-(N-s)(N-s-1)}\sqrt{N\left(N+1\right)-(N-s-|l|)(N-s-|l|-1)},$
$d^{(2)}_{(l,s)}$ $=\frac{1}{8N(2N-1)^{2}}\sqrt{(2N-s)(2N-s-1)(s+1)(s+2)}\sqrt{(2N-s-|l|)(2N-s-|l|-1)(s+|l|+1)(s+|l|+2)},$
while $d^{(k\geq 3)}_{(l,s)}=0$. Using Eq. (S40) and Eq. (S25), we can construct the complete projection of $\left(\mathcal{E}\!+\!\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}\right)$ to the degenerate subspace of the $l$th layer.
Figure S2: The real part of PXP Liouvillian spectrum $\mathrm{Re}(\lambda^{\prime}/\gamma)$, as the dephasing rate $\gamma$ is varied. The results are calculated by (a) including the linear ($\hat{\tau}^{z}_{b}$) and the quadratic ($|0\rangle\langle 0|_{b}=\hat{I}-(\hat{\tau}^{z}_{b})^{2}$) terms, (b) only the linear term, (c) only the quadratic them. The vertical dashed lines indicate the spectral transition points. The system size is chosen by $L=10$.
The results of perturbation theory of the PXP model are plotted in Fig. (S2) against the numerical results for the same system size. While the full result is in good agreement with numerics, we see that dropping the quadratic term in Eq. (S27), shown in panel (b), already captures the spectral transitions. In fact, in this case we see the transitions become sharply defined, reminiscent of the CL model in the main text. This is not a coincidence since the perturbed PXP Liouvillian without the quadratic term obeys the $\mathbb{PT}$ symmetry, which is generated by the superoperator $\mathcal{T}_{-}$,
$\mathcal{T}_{-}=\sqrt{\pm 1}\Big[\Big(\prod_{b}\hat{\tau}_{b}^{x}\Big)\otimes\hat{I}\Big]\mathcal{K},\quad\quad\mathcal{T}_{-}\mathcal{L}\mathcal{T}_{-}^{-1}=-\mathcal{L},$ (S41)
where $\mathcal{K}$ is the complex-conjugation superoperator. Conversely, we do not observe any spectral transitions if we only consider the quadratic term and discard the linear term – see panel (c) in Fig. (S2). This suggests that for models like PXP, the spectral transition experienced by QMBS Liouvillian eigenvalues is not due to spontaneous breaking of an exact symmetry generated by $\mathcal{T}_{-}$.
S1.3 A3. Scaling behavior
@@
From the perspective of perturbation theory, the spectral transition phenomenon originates from the competition between dissipation $\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}$ and the self-energy $\mathcal{E}$, as summarized in Fig. S1. We can estimate the typical Liouvillian spectral transition point:
@@
where $\|\hat{A}\|=\sqrt{\lambda_{\mathrm{max}}(\hat{A}^{\dagger}\hat{A})}$ refers to the matrix norm of $\hat{A}$. The dissipation superoperator scales with $N$, $\|\mathcal{D}^{\prime}\|=\sum_{j=1}^{N}\|\hat{\sigma}_{j}^{z}\otimes\hat{\sigma}_{j}^{z}\|\sim N$. To estimate the scaling of self-energy for large $N$, recall the self-consistency condition, Eq. (S13), and focus on the CL model, with $\hat{S}_{0}$ in Eq. (S6). The maximal energy is $NJ$, so we have $\|E\hat{I}-\hat{S}_{0}\|$ scaling as $N$. According to Eq. (S14),
$\big\|\hat{\Gamma}\big\|=\sqrt{\max{\bigg(\sum_{j=1}^{N-1}J_{\mathrm{x},j}\hat{P}^{j,j+1}_{p_{\mathrm{S}}}\bigg)}}.$ (S43)
The possible value of $\hat{P}^{j,j+1}_{p_{\mathrm{S}}}$ is 0 and $1$, so $\|\hat{\Gamma}\|=\sqrt{\sum_{j=1}^{N-1}J_{\mathrm{x},j}}\sim\sqrt{N}$, where we assume that all the couplings $J_{\mathrm{x},j}$ are at the same order of magnitude. Approximately, from the self-consistence equation, we assume that $\|\hat{\Sigma}\|$ increase slower than $\|E\hat{I}-\hat{S}_{0}\|$, and we have